A service station checks Mr. Gittleboro's radiator and finds it contains only 30% antifreeze. If the radiator holds 10 quarts and is full, how much be drained off and replaced with pure antifreeze in order to bring it up to a required 50% anti freeze?
Answer by mananth(16946)

Assume we have to removae x quarts of antifreeze from the radiator and add 100% one
Antifreeze
percent ---------------- quantity
Antifreeze pure 100.00% ---------------- x quarts
Antifreeze I 30.00% ------ 10 - x quarts
Mixture 50.00% ---------------- 10
10
100.00% x + 30.00% ( 10 - x ) = 50.00% * 10
100 x + 30 ( 10 - x ) = 500
100 x + 300 - 30 x = 500
100 x - 30 x = 500 - -300
70 x = 200
/ 70
x = 2.86 quarts 100.00% Antifreeze pure
7.14 quarts 30.00% Antifreeze I

m.ananth@hotmail.ca